# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

### v1

@@ IV-D4 Qualitative Results
The LAM comparisons in Fig. 13, based on Figs. 10 and 12, demonstrated that GDSR benefitted from the superior long-range modeling capability of RGEG, enabling it to leverage more pixels during the SR process. Additionally, the strong detail extraction capability of RDEG allowed GDSR to utilize pixels correctly for reconstruction. Specifically, for the RSSRD-QH_CDM dataset, when large-scale and global information was present, CNN- and Transformer-based SR networks struggled to extract sufficient information, and while the Mamba-based SR network activated a large number of pixels, they failed to use them correctly for reconstruction. In contrast, thanks to the effective integration of RGEG and RDEG, the GDSR-generated results significantly outperformed other SR models. Moreover, GDSR surpassed MambaIR in terms of a diffusion index by 25.6193, indicating that by incorporating an RWKV with a larger receptive field into RGEG, GDSR could better capture global information in images with more such features, leading to a higher diffusion index. These visual comparisons further validated the effectiveness of GDSR in capturing and reconstructing details in remote sensing images, highlighting the powerful pixel utilization capability brought by our global-detail dual-branch structure.
@@ IV-D5 Real-World Image Testing on the Sanjiangyuan Region
To further evaluate the model’s performance in real-world scenarios, we tested it on remote sensing images from real-world environments. As shown in Fig. 14, the image depicts an alpine meadow grassland type, with surface features such as cracks, scattered rocks, and numerous small ”mouse holes.” These mouse holes are depressions or openings on the ground, and the distribution of rodents is significant for studying the primary grassland type, the alpine meadow, in Qinghai Province. Their presence can impact the biodiversity, productivity, and ecological function of the alpine meadows. However, due to the limitations in drone flight conditions and resolution, the raw images suffer from motion blur and loss of details. These issues hinder accurate observation of key features such as surface textures and the depth of the mouse holes. By applying SR techniques, our GDSR effectively enhanced the clarity of these fine details, sharpened the edges around the mouse holes, and restored the surface texture. This improvement is crucial for accurate environmental analysis, as the restored landscape provides more precise data for studying biodiversity conservation and the enhancement of ecological functions in the alpine meadow ecosystem of Qinghai Province.
V Conclusion
@@
In this study, we first introduce the Receptance Weighted Key Value for remote sensing image super-resolution. GDSR effectively models the global dependencies of large-scale remote sensing images while incorporating local details, all with linear complexity and significantly improved inference speed. Specifically, we design an efficient and effective Global-Detail dual-branch architecture that correctly activates more pixels spatially for image reconstruction. To combine global and local representations, we propose GDRM, which adaptively adjusts the features of the dual branches and fuses them. Additionally, we introduce Wavelet Loss to help the model better capture high-frequency details in remote sensing image super-resolution tasks. Extensive quantitative and qualitative experiments on the AID, AID_CDM, RSSRD-QH, and RSSRD-QH_CDM benchmarks demonstrate that our GDSR outperforms state-of-the-art CNN-, Transformer-, and Mamba-based SR models in remote sensing image super-resolution tasks. The proposed Wavelet Loss also enhances the performance of various models in remote sensing image super-resolution tasks.
This study focuses solely on ×3 SR, limiting the flexibility to explore SR at different scales. Experiments show that excessive high-frequency components introduced by Wavelet Loss can lead to a decline in perceptual quality. In future work, we plan to extend our GDSR to include more scaling factors, further demonstrating its robustness and effectiveness. Additionally, we aim to optimize Wavelet Loss to improve its perceptual enhancement.

### v4

@@ IV-D4 Qualitative Results
The LAM comparisons in Fig. 13, based on Figs. 10 and 12, demonstrated that GDSR benefited from the superior long-range modeling capability of RGEG, enabling it to leverage more pixels during the SR process. Additionally, the strong detail extraction capability of RDEG allowed GDSR to utilize pixels correctly for reconstruction. Specifically, for the RSSRD-QH_CDM dataset, when large-scale and global information was present, CNN- and Transformer-based SR networks struggled to extract sufficient information, and while the Mamba-based SR network activated a large number of pixels, it failed to use them correctly for reconstruction. In contrast, thanks to the effective integration of RGEG and RDEG, the GDSR-generated results significantly outperformed other SR models. Moreover, GDSR surpassed MambaIR in terms of a diffusion index by 25.6193, indicating that by incorporating an RWKV with a larger receptive field into RGEG, GDSR could better capture global information in images with more such features, leading to a higher diffusion index. These visual comparisons further validated the effectiveness of GDSR in capturing and reconstructing details in remote sensing images, highlighting the powerful pixel utilization capability brought by our global-detail dual-branch structure.
@@ IV-D5 Real-World Image Testing on the Sanjiangyuan Region
To further evaluate the model’s performance in real-world scenarios, we tested it on RSIs from real-world environments. As shown in Fig. 14, the image depicts an alpine meadow grassland type, with surface features such as cracks, scattered rocks, and numerous small ”mouse holes.” These mouse holes are depressions or openings on the ground, and the distribution of rodents is significant for studying the primary grassland type, the alpine meadow, in Qinghai Province. Their presence can impact the biodiversity, productivity, and ecological function of the alpine meadows. However, due to the limitations in drone flight conditions and resolution, the raw images suffer from motion blur and loss of details. These issues hinder accurate observation of key features such as surface textures and the depth of the mouse holes. By applying SR techniques, our GDSR effectively enhanced the clarity of these fine details, sharpened the edges around the mouse holes, and restored the surface texture. This improvement is crucial for accurate environmental analysis, as the restored landscape provides more precise data for studying biodiversity conservation and the enhancement of ecological functions in the alpine meadow ecosystem of Qinghai Province.
V Discussion
The proposed method demonstrated consistent SR performance across diverse remote sensing scenarios, validated on the AID (30 land-use categories) [44], UCMerced (21-class geospatial scenes) [45], and custom RSSRD-QH (Qinghai Plateau ecology) datasets. By enhancing texture fidelity in complex environments such as fragmented agricultural landscapes, high-density urban settlements, and ecologically sensitive alpine regions, this approach supports critical applications requiring precise feature extraction. Improved road network reconstruction in severely degraded AID samples aids emergency routing during disasters, while sharpened landscape details in RSSRD-QH enable accurate environmental monitoring. The method’s adaptability to both global benchmarks and region-specific challenges positions it as a practical tool for revitalizing historical mid-resolution archives into high-precision analysis-ready data.
Additionally, we tested GDSR’s robustness against challenging, remote sensing-specific artifacts. While our initial CDM followed established blind SR methods [49, 50, 51, 52], real-world RSIs often contain severe sensor-induced stripe noise. To validate our method’s adaptability, we augmented the CDM with this artifact and retrained the GDSR model. As demonstrated in Fig. 15, the results confirmed GDSR’s strong adaptive capabilities. The retrained model successfully identified and eliminated severe stripe noise in images, leading to quantitative and visual improvements. This positive result paves the way for future work to incorporate other challenging degradations, such as atmospheric scattering, further enhancing the model’s practical utility.
VI Conclusion
@@
In this study, we first introduce the RWKV for RSISR. GDSR effectively models the global dependencies of large-scale RSI while incorporating local details, all with linear complexity and significantly improved inference speed. Specifically, we design an efficient and effective Global-Detail dual-branch architecture that correctly activates more pixels spatially for SR. To combine global and local representations, we propose GDRM, which adaptively adjusts the features of the dual branches and fuses them. Additionally, we introduce Dual-Group Multi-Scale Wavelet Loss to help the model better reconstruct high-frequency details. Extensive quantitative and qualitative experiments on the AID, UCMerced, and RSSRD-QH benchmark datasets using two degradation methods demonstrate that our GDSR outperforms state-of-the-art CNN-, Transformer-, and Mamba-based SR models in RSI-SR tasks.
Nevertheless, our GDSR does exhibit some shortcomings. First, the use of simple stacked residual convolutions for detail extraction introduces a high parameter count, limiting model efficiency. Future work will explore more efficient architectures specifically tailored for RSIs. Second, the current validation is limited to RGB RSIs, lacking evaluation on hyperspectral SR scenarios where balancing spatial enhancement with spectral fidelity is paramount. Extending the method to hyperspectral data and rigorously assessing its spectral accuracy will be a key future direction. Moreover, we consider investigating the model’s applicability to multi-temporal or cross-sensor RSIs, addressing challenges like temporal consistency and domain adaptation to further enhance its practical utility.